OpenStudy (anonymous):

A shipping company uses a formula to determine the cost for shipping a package: c = 2.79 + 0.38p, where c is the cost of shipping and p is the number of pounds. What is the cost of shipping a package that weighs 8 pounds? A. $2.79 B. $3.04 C. $5.21 D. $5.83

OpenStudy (realitywillslapyou):

Here, we have c, which is the cost of shipping and it is expressed by formula c = 2.79 + 0.38*p. If we know that the weight of the package is 8 pound, we just replace p in formula with 8: c = 2.79 + 0.38 * p c = 2.79 + 0.38 * 8 c = 2.79 + 3.04 c = 5.83 So, the correct answer is d. $5.83 I hope that this helps! My Source: http://www.homeworkmarket.com/content/shipping-company-uss-formula-determine-cost-shipping-package-c279038p-where-c-cost